OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x - y = 3 7x - 9y = -2

OpenStudy (agent47):

4x-3=y 7x-9(4x-3)=-2 7x-36x+27=-2 -29x=-2-27 -29x=-29 x=1 4*1-3=y 4-3=y y=1

OpenStudy (anonymous):

x=1,y=1